OpenStudy (anonymous):

I need some homework help? Use the quadratic formula to solve 10x2 - 30x – 100 = 0. A. x = -2 or x = 5 B. x = -5 or x = 2 C. x = 2 or x = -4 D. x = -4 or x = -2

OpenStudy (solomonzelman):

Do you know the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that's why im asking what the answer is

OpenStudy (solomonzelman):

\[\huge\color{blue}{ x=~~\frac{-b±\sqrt{b^2-4(a)(c)} }{2a} } \]

OpenStudy (solomonzelman):

Are you familiar with this formula in blue ?

OpenStudy (anonymous):

I've seen it before, I'm just not sure how to do it

OpenStudy (solomonzelman):

alright, the a b c there are numbers. the standard form of the equation is \[\huge\color{blue}{ ax^2+bx+c=0 } \] compare your equation to this form and tell me what the "a" b" and "c" in your equation are

OpenStudy (solomonzelman):

\[ \huge\color{blue}{ \color{red} { 10 } x^2\color{red} { -30 } x\color{red} { -100 } =0 } \] can you tell me the A B C now ?

OpenStudy (anonymous):

would a be 10 and 30 be b and 100 be c? I'm confused still

OpenStudy (solomonzelman):

Yeah, exactly ! now you need to plug them in (that's all) !

OpenStudy (solomonzelman):

\[\color{purple}{ x=~~\frac{-(-30)±\sqrt{(-30)^2-4(10)(-100)} }{2(10)} } \] \[\color{purple}{ x=~~\frac{+30±\sqrt{(-30)^2-(-4000)} }{2(10)} } \] \[\color{purple}{ x=~~\frac{+30±\sqrt{900+4000} }{2(10)} } \] \[\color{purple}{ x=~~\frac{+30±\sqrt{4900} }{20} } \] \[\color{purple}{ x=~~\frac{+30±70 }{20} } \] \[\color{brown}{ x=~5~~~~~or~~~~~ -2 } \]

OpenStudy (solomonzelman):

See how I am doing this ?

OpenStudy (anonymous):

Yes i see i think. It's just the + or - symbols really confuse me..I never learned how to do that

OpenStudy (solomonzelman):

this is just because of a square root. b/c you know that \[\sqrt{49}=+7~~~~or~~~-7\]

OpenStudy (anonymous):

But from what I got from doing all of that I got x = -2 or x = 5

OpenStudy (anonymous):

and that makes much more sense now

OpenStudy (solomonzelman):

Yup ! Nice to hear that. Good luck !

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

np